# Jamais Vu: Exposing the Generalization Gap in Supervised Semantic Correspondence

### v1

@@ Abstract
Semantic correspondence (SC) aims to establish semantically meaningful matches across different instances of an object category. We illustrate how recent supervised SC methods remain limited in their ability to generalize beyond sparsely annotated training keypoints, effectively acting as keypoint detectors. To address this, we propose a novel approach for learning dense correspondences by lifting 2D keypoints into a canonical 3D space using monocular depth estimation. Our method constructs a continuous canonical manifold that captures object geometry without requiring explicit 3D supervision or camera annotations. Additionally, we introduce SPair-U, an extension of SPair-71k with novel keypoint annotations, to better assess generalization. Experiments not only demonstrate that our model significantly outperforms supervised baselines on unseen keypoints, highlighting its effectiveness in learning robust correspondences, but that unsupervised baselines outperform supervised counterparts when generalized across different datasets.
@@ 1 Introduction
The task of semantic correspondence (SC) estimation involves identifying semantically matching regions across different instances of the same object category. It remains a challenging problem, as it requires recovering fine-grained details while maintaining robustness against variations in appearance and shape. Recent advances in large-scale vision models, particularly self-supervised transformers caron2021emerging ; oquab2023dinov2 and generative diffusion models rombach2022high , have led to notable improvements in SC. When employed as backbones, these models have achieved over 20% gains in accuracy on the SPair-71k benchmark min2019spair . However, despite these advances, recent studies have highlighted that these powerful representations often struggle to disambiguate symmetric object parts due to their visual similarity zhang2023telling ; mariotti2024improving .
@@
SC methods can be broadly categorized into two groups in terms of supervision: unsupervised models, which do not require correspondence annotations during training amir2021deep ; aygun2022demystifying ; taleof2feats ; mariotti2024improving , and supervised models cho2021cats ; huang2022learning ; zhang2023telling , which are trained on annotated correspondences. As expected, supervised models generally achieve higher performances when using the same backbone and same training set as unsupervised models. However, a key limitation of current benchmarks is that evaluation is typically performed on the same set of keypoints used for training, potentially inflating perceived generalization. As illustrated in Fig. 1, the performance of supervised models drops significantly when evaluated on unseen keypoints, while unsupervised models maintain their performance.
@@
Figure 1: Illustration of the generalization gap on unseen keypoints. (Left) Top row: when evaluated on known keypoints, both our model and Geo-SC zhang2023telling perform well while the unsupervised DINOv2+SD taleof2feats struggles to correctly disambiguate the legs of the horse. Bottom row: when presented with keypoints unseen at training time, both our model and DINO+SD predict noisy but reasonable correspondence, while Geo-SC predictions noticeably degrade. (Right) Even though it obtains strong performance on known keypoints, Geo-SC performs worse than its unsupervised counterpart on our new benchmark of unseen keypoints. In comparison, our model still achieves competitive results.
@@
Building dense correspondences are key to fine-grained object understanding and improving robustness in various recognition tasks, and essential to many applications including texture transfer ofri2023neural and robotic manipulation tsagkas2024click . In this paper, we examine the performance of state-of-the-art SC models when evaluated on points that lie outside the set of annotated training keypoints. Under these conditions, we observe that supervised pipelines often underperform their unsupervised counterparts, effectively reducing their function to that of ‘sparse keypoint detectors’. We attribute this limitation to two main factors: (i) the sparsity of supervision, which typically focuses on a limited number of keypoints and (ii) the lack of evaluation on unseen points, which tends to favor models that bias their predictions toward the nearest seen annotations.
@@
We argue that an ideal SC method should be capable of matching arbitrary points, akin to the objectives in classical dense correspondence tasks liu2010sift . To move beyond sparse supervision, we propose a learning framework that predicts dense feature maps and supervises them using geometry from an off-the-shelf depth estimator, thereby enabling training on richer and more spatially diverse cues. Few unsupervised SC methods thewlis2017unsupervised ; mariotti2024improving have leveraged 3D geometry to learn dense correspondences through mapping object pixels to a spherical coordinate system where each coordinate corresponds to a different characteristic point of the object. However, this approach requires estimating the object shape and viewpoint from a collection of 2D images, which limits to the applicability of such models to synthetically generated datasets thewlis2017unsupervised or relying on camera viewpoint supervision mariotti2024improving .
@@
We propose a new approach that leverages existing 2D keypoint annotations and estimated 3D geometry to learn dense correspondences without requiring any direct 3D supervision. We build on the idea of learning a canonical representation of the object category, which is invariant to the object instance, viewpoint, and pose. We achieve this by lifting the 2D keypoints to 3D using a monocular depth model, aligning them with a set of canonical keypoints that are shared across all instances of the object category. Finally, by interpolating between them, we learn a continuous canonical manifold, that captures the underlying 3D shape of the object and incorporates geometric constraints into learning more effective and general feature representations. We also introduce a new dataset for SC, SPair-U which extends the original SPair-71k test annotations with a set of new keypoints, allowing us to evaluate the generalization of SC models to unseen keypoints. We show that supervised SC models trained on the original SPair-71k dataset typically fail to generalize well to unseen keypoints, while our method is able to learn a more general representation that can be applied to unseen keypoints.
@@ 2 Related Work
Supervised methods. Supervised approaches rely on the availability of datasets with annotated keypoints such as CUB wah2011cub , PF-PASCAL ham2017proposal , and SPair-71k min2019spair to learn corresponding points across instances of the same object class depicted in different images. This is typically using contrastive objectives minimizing distance between features coming from the same keypoints while pushing other features away han2017scnet ; tang2023emergent ; taleof2feats ; zhang2023telling . Flow-based approaches, that compute 4D correlations maps between source and target images have also been explored cho2021cats ; huang2022learning . To obtain stronger descriptors, it is also common to aggregate features from multiple network layers to form hypercolumns min2020learning ; amir2021deep ; zhang2023telling . Current state-of-the-art supervised methods forego training from scratch and instead typically use a large pretrained vision model as a backbone, the most popular options being DINOv2 oquab2023dinov2 and Stable Diffusion rombach2022high . While effective at matching instances of keypoints of the same type that have been observed during training, in our experiments we demonstrate that current supervised methods have a tendency to overfit to the set of keypoints observed during training and struggle to generalize to previously unseen keypoints (see Fig. 1 for an example).
@@
Unsupervised / weakly-supervised methods. Methods that do not use correspondence supervision during training range from completely unsupervised approaches thewlis2017unsupervised ; thewlis2019unsupervised ; aygun2022demystifying ; amir2021deep ; taleof2feats , zero-shot methods that only use test-time information about the relationships between keypoints ofri2023neural ; zhang2023telling , dense methods that directly impose structure on the correspondence field cheng2024zero , and weakly supervised methods that use extra labels like segmentation masks or camera pose kanazawa2016warpnet ; mariotti2024improving ; barel2024spacejam . Earlier unsupervised methods typically use self-supervised objectives that make use of synthetic deformations/augmentations of the same image thewlis2017unsupervised or by using cycle consistencies thewlis2019unsupervised to provide pseudo ground truth correspondence. Later it was observed that large pretrained vision models naturally posses features that are very strong for SC, despite not being trained on this task explicitly. As a result, more recent unsupervised methods tend to not train their own backbones from scratch and instead explore ways to aggregate amir2021deep ; taleof2feats , or align gupta2023asic ; mariotti2024improving these features across images.
@@
Geometry-aware methods. Inspired by the classic correspondence setting in vision that relies on geometric constraints to match the same 3D locations across views davison2007monoslam ; schonberger2016structure , utilizing geometry cues is an effective way to learn SC. For SC, the underlying assumption is that different instances from the same object category share a similar spatial structure. Flow and rigidity constraints are often used in tracking wang2023tracking ; xiao2024spatialtracker and unsupervised SC gupta2023asic ; barel2024spacejam . Recent studies have shown that ambiguities caused by symmetric objects are a major source of errors in SC zhang2023telling ; mariotti2024improving . One potential way to mitigate these errors is to develop 3D-aware methods. Initially, this has been explored by building correspondence across images by matching points along the surface of objects to 3D meshes zhou2016learning ; kulkarni2019canonical ; neverova2020continuous , but the requirement for meshes greatly limits the applicability of these approaches.
More recently, methods have been proposed to learn 3D shape from image collections monnier2022share ; wu2023magicpony ; aygun2023saor . However, these methods require solving multiple problems at once, i.e., estimating object shape, deformations, camera pose, and are therefore limited to specific types of object categories and tend to break easily when applied to more complex shapes. Recent advances in monocular depth ranftl2020towards ; bhat2023zoedepth ; ke2024repurposing ; yang2024depth ; fu2024geowizard ; wang2024moge and geometry prediction wang2024dust3r allows for reliable geometry estimation from a single image, which can be leveraged for imparting 3D-awareness in into SC methods.
Concurrent work wandel2025semalign3d , also proposes to geometrically align images in 3D using depth maps to build category prototypes. However, it is designed around a test-time alignment of the 3D prototype to the test image, which requires segmentation masks, and lacks the ability to generalize to new categories, as no prototype for them would exist in absence of training samples.
@@ 3.1 Overview
Figure 2: Overview of our approach. We extract segmentation masks and depths maps for training images and backproject object points to produce the posed point clouds $\mathcal{X}^{b}$. We predict dense features with $\Phi$ and match them against our jointly learned sparse category prototype $(\mathcal{P},\mathcal{Z})$ to produce the canonical point clouds $\mathcal{X}^{c}$. The local geometric alignment between the two provides supervision for updating $\Phi$.
Let $I\in\mathbb{R}^{H\times W\times 3}$ be an RGB image of the object, defined over the image domain $\Lambda\in\mathbb{R}^{2}$, a lattice of size $H\times W$. Our objective is to learn a function $\Phi(I,\bm{u})\rightarrow\bm{z}$, which maps each pixel coordinate $\bm{u}\in\Lambda$ to a descriptor $\bm{z}\in\mathbb{R}^{M}$. The descriptor $\bm{z}$ should be semantically consistent (e.g., left eye of a cat), meaningfully aligned across different images of objects from the same category, and be invariant to changes in pose and shape. Once $\Phi$ is learned, semantic correspondence can be established by finding, for a pixel $\bm{u}$ in image $I$, the most similar pixel $\bm{u}^{\prime}$ in another image $I^{\prime}$ of the same object category. This is done by querying nearest-neighbor matching in descriptor space according to distance $d$, typically the cosine distance, i.e., $d(\bm{a},\bm{b})=1-\langle\bm{a},\bm{b}\rangle/(||\bm{a}||_{2}||\bm{b}||_{2})$:
@@
In the standard supervised SC task, we are given a training set, $\{(I^{(1)},\mathcal{K}^{(1)}),\dots,(I^{(N)},\mathcal{K}^{(N)})\}$ where each image $I^{(n)}$ is annotated with a sparse set of semantic keypoints $\mathcal{K}=\{{\bm{k}}_{1},\dots,{\bm{k}}_{\lvert\mathcal{K}\rvert}\}$ where ${\bm{k}}\in\Lambda$. A common strategy, adopted in recent works huang2022learning ; cho2021cats ; cho2022cats++ ; luo2023diffusion ; taleof2feats ; zhang2023telling , is to learn a descriptor function $\Phi(I,\bm{u})$ that produces a local descriptor $\bm{z}$ for each pixel $\bm{u}$ in $I$ such that the descriptors of corresponding keypoints in paired images are close in feature space. While this sparse keypoint supervision helps the model learn semantically meaningful descriptors for the annotated keypoints, it does not guarantee consistency for unlabeled regions of the object.
A promising direction to address this limitation is to incorporate 3D geometry by assigning each pixel a coordinate in an object centric reference frame. Prior works thewlis2017unsupervised ; mariotti2024improving explores this idea by projecting object surfaces onto a spherical coordinate system, with each coordinate on the sphere corresponding to a different characteristic point of the object. However, this requires inferring both object shape and the viewpoint from a collection of 2D images, a highly ill-posed problem, which requires generating pairs through synthetic warps resulting unrealistic shapes thewlis2017unsupervised or relies on viewpoint supervision which can be challenging to predict automatically. Next we show how to combine sparse keypoints annotations with 3D geometry cues to learn dense, semantically consistent descriptors for every pixel in an image. An overview of our approach is shown in Fig. 2.
@@ 3.2 Canonical Representation Learning
Similar to thewlis2017unsupervised ; mariotti2024improving , we aim to learn a 3D canonical representation for each object category, along with a function $\varphi(I,\bm{u})\rightarrow\bm{x}\in\mathbb{R}^{3}$ that maps a pixel $\bm{u}$ in image $I$ to its 3D coordinates in the canonical object-centric coordinate system. Unlike the spherical representations in prior work thewlis2017unsupervised ; mariotti2024improving , we do not impose any topological constraints on the canonical representation (e.g.,, enforcing a spherical surface). We parameterized the canonical representation by a set of 3D keypoints $\mathcal{P}=\{\bm{p}_{1},\dots,\bm{p}_{\lvert\mathcal{K}\rvert}\}$, where each $\bm{p}_{i}$ corresponds to a labeled keypoint ${\bm{k}}_{i}$ in $I$. Crucially, $\mathcal{P}$ is shared across all instances of a category, and is invariant to object instance, viewpoint, and pose, unlike $\mathcal{K}$, ensuring the canonicity of the representation.
To compute $\mathcal{P}$, we first compute the 3D coordinates of each keypoint ${\bm{k}}$ in image $I$ using a monocular depth model $\Psi(I,\bm{k})\rightarrow\mathbb{R}^{+}$ and then backproject it by using estimated camera intrinsics $\mathbf{A}\in\mathbb{R}^{3\times 3}$ as follows:
@@
To align $\bar{\mathcal{K}}$ with $\mathcal{P}$, we compute a rigid transformation, comprising rotation $\mathbf{R}\in SO(3)$ and translation $\mathbf{T}\in\mathbb{R}^{3}$, and scale $s\in\mathbb{R}^{+}$ such that $\mathbf{M}=s[\mathbf{R}|\mathbf{T}]\in\mathbb{R}^{3\times 4}$. We optimize the canonical keypoints $\mathcal{P}$ by minimizing their alignment error across the training set using a nested optimization:
$\min_{\bm{p}_{i}}\sum_{n=1}^{N}||\bm{p}_{i}-\hat{\bm{M}}^{(n)}\bar{\bm{k}}_{i}^{(n)}||_{1}\quad\text{ where }\quad\hat{\bm{M}}^{(n)}=\arg\min_{\bm{M}}\sum_{i=1}^{\lvert\mathcal{K}\rvert}||\bm{p}_{i}-\bm{M}\bar{\bm{k}}_{i}^{(n)}||_{2}.$ (3)
In the inner loop, we use Kabsch’s algorithm kabsch1976solution to efficiently compute the optimal rigid alignment $\hat{\bm{M}}^{(n)}$ between the canonical and posed coordinates. Even though alignments rely on only a sparse subset of visible keypoints per-image, we find this sufficient to recover a globally consistent arrangement of the $\mathcal{P}$.
@@
with cosine similarity sim(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) and learned temperature parameter $\tau$. This objective encourages $\Phi$ to produce distinctive and semantically consistent features across object instances for each keypoint $\bm{k}$.

### v2

@@ Abstract
The goal of semantic correspondence (SC) estimation is to establish semantically meaningful matches across different instances of an object category. In this work, we illustrate how recent supervised SC methods generalize poorly beyond the annotated keypoints seen during training, thus effectively acting as keypoint detectors. To address this, we propose a new approach for learning dense correspondences by lifting 2D keypoints into a canonical 3D space using monocular depth estimation. Our method constructs a continuous canonical manifold that captures object geometry without requiring explicit 3D supervision or camera annotations. Additionally, we introduce SPair-U, an extension of SPair-71k with novel keypoint annotations, to better assess generalization. Experiments not only demonstrate that our model significantly outperforms supervised baselines on unseen keypoints, highlighting its effectiveness in learning robust correspondences, but that unsupervised baselines outperform supervised counterparts when evaluated across different datasets.
@@ 1 Introduction
Semantic correspondence (SC) estimation involves identifying semantically matching regions in images across different instances of the same object category. It remains a challenging problem, as it requires recovering fine-grained details while maintaining robustness against variations in object appearance, shape, and viewing conditions. Recent advances in large-scale vision models, particularly self-supervised transformers caron2021emerging; oquab2023dinov2 and generative diffusion models rombach2022high, have led to notable improvements in SC. When employed as backbones, these models have achieved over 20% gains in accuracy on the SPair-71k benchmark min2019spair. However, despite these advances, recent studies have highlighted that these powerful representations often struggle to disambiguate symmetric object parts due to their visual similarity zhang2023telling; mariotti2024improving.
@@
SC methods can be broadly categorized into two groups in terms of supervision: unsupervised models, which do not require correspondence annotations during training amir2021deep; aygun2022demystifying; taleof2feats; mariotti2024improving, and supervised models cho2021cats; huang2022learning; zhang2023telling, which are trained on manually annotated correspondences. As expected, supervised models generally achieve higher performances when using the same backbone and same training set as unsupervised models. However, a key limitation of current benchmarks is that evaluation is typically performed on the same set of keypoints used for training, potentially inflating perceived generalization. As illustrated in Fig.˜1, the performance of supervised models drops significantly when evaluated on unseen keypoints, while unsupervised models maintain their performance.
@@
Figure 1: Illustration of the generalization gap on unseen keypoints. (Left) Top row: when evaluated on known keypoints, both our model and Geo-SC zhang2023telling perform well, while the unsupervised DINOv2+SD taleof2feats struggles to correctly disambiguate the legs of the horse. Bottom row: when presented with keypoints unseen at training time, both our model and DINO+SD predict noisy but reasonable correspondence, while Geo-SC predictions noticeably degrade. (Right) Even though it obtains strong performance on known keypoints, Geo-SC performs worse than its unsupervised counterpart on our new benchmark of unseen keypoints. In comparison, our model still achieves competitive results.
@@
Building dense correspondences are key to fine-grained object understanding and improving robustness in various recognition tasks, and essential to many applications including texture transfer ofri2023neural and robotic manipulation tsagkas2024click. In this work, we examine the performance of state-of-the-art SC models when evaluated on points that lie outside the set of annotated training keypoints. Under these conditions, we observe that supervised pipelines often underperform their unsupervised counterparts, effectively reducing their function to that of ‘sparse keypoint detectors’. We attribute this limitation to two main factors: (i) the sparsity of supervision, which typically focuses on a limited number of keypoints and (ii) the lack of evaluation on unseen points, which tends to favor models that bias their predictions toward the nearest seen annotations.
@@
We argue that an ideal SC method should be capable of matching arbitrary points, akin to the objectives in classical dense correspondence tasks liu2010sift. To move beyond sparse supervision, we propose a learning framework that predicts dense feature maps and supervises them using geometry from an off-the-shelf depth estimator, thereby enabling training on richer and more spatially diverse cues. Some unsupervised SC methods thewlis2017unsupervised; mariotti2024improving have leveraged 3D geometry to learn dense correspondences through mapping object pixels to a spherical coordinate system where each coordinate corresponds to a different characteristic point of the object. However, this approach requires estimating the object shape and viewpoint from a collection of 2D images, which limits the applicability of such models to synthetically generated datasets thewlis2017unsupervised and they can require additional camera viewpoint supervision mariotti2024improving.
@@
We propose a new approach that leverages existing 2D keypoint annotations and estimated 3D geometry to learn dense correspondences. We build on the idea of learning a canonical representation of the object category, which is invariant to the object instance, viewpoint, and pose. We achieve this by lifting the 2D keypoints to 3D using a monocular depth model, aligning them with a set of canonical keypoints that are shared across all instances of the object category. Finally, by interpolating between them, we learn a continuous canonical manifold, that captures the underlying 3D shape of the object and incorporates geometric constraints into learning more effective and general feature representations. We also introduce a new dataset for SC estimation, SPair-U, which extends the original SPair-71k test annotations with a set of new keypoints, allowing us to evaluate the generalization of SC models on unseen keypoints. We show that supervised SC models trained on the original SPair-71k dataset typically fail to generalize well to unseen keypoints, while our method is able to learn a more general representation that can be applied to unseen keypoints.
@@ 2 Related Work
Supervised methods. Supervised approaches rely on the availability of datasets with annotated keypoints such as CUB wah2011cub, PF-PASCAL ham2017proposal, and SPair-71k min2019spair to learn corresponding points across instances of the same object class depicted in different images. This is typically using contrastive objectives minimizing distance between features coming from the same keypoints while pushing other features away han2017scnet; tang2023emergent; taleof2feats; zhang2023telling. A more computationally intensive option is to compute dense 4D correlations maps between each source and target locations cho2021cats; min2021convolutional; huang2022learning; kim2022transformatcher. To obtain stronger descriptors, it is also common to aggregate features from multiple network layers to form hypercolumns min2020learning; amir2021deep; zhang2023telling. Current state-of-the-art supervised methods forego training from scratch and instead typically use a large pretrained vision model as a backbone, the most popular options being DINOv2 oquab2023dinov2 and Stable Diffusion rombach2022high. While effective at matching instances of keypoints of the same type that have been observed during training, in our experiments we demonstrate that current supervised methods have a tendency to overfit to the set of keypoints observed during training and struggle to generalize to previously unseen keypoints (see Fig.˜1 for an example).
@@
Unsupervised / weakly-supervised methods. Methods that do not use correspondence supervision during training range from unsupervised approaches using general-purpose backbones amir2021deep; taleof2feats, very weakly supervised methods that only assume an curated training set without labels, e.g., images of a single category thewlis2017unsupervised; thewlis2019unsupervised; aygun2022demystifying, zero-shot methods that only use test-time information about the relationships between keypoints ofri2023neural; zhang2023telling, dense methods that directly impose structure on the correspondence field cheng2024zero, and weakly supervised methods that use extra labels like segmentation masks or camera pose kanazawa2016warpnet; mariotti2024improving; barel2024spacejam. Earlier unsupervised methods typically use self-supervised objectives that make use of synthetic deformations/augmentations of the same image thewlis2017unsupervised or by using cycle consistencies thewlis2019unsupervised; truong2021warp; truong2022probabilistic to provide pseudo ground truth correspondence. Later it was observed that large pretrained vision models naturally posses features that are very strong for SC, despite not being trained on this task explicitly. As a result, more recent unsupervised methods tend to not train their own backbones from scratch and instead explore ways to aggregate amir2021deep; taleof2feats, or align gupta2023asic; mariotti2024improving these features across images.
@@
Geometry-aware methods. Inspired by the classic correspondence setting in vision that relies on geometric constraints to match the same 3D locations across views davison2007monoslam; schonberger2016structure, utilizing geometry cues is an effective way to learn SC. For SC, the underlying assumption is that different instances from the same object category share a similar spatial structure. Flow and rigidity constraints are often used in tracking wang2023tracking; xiao2024spatialtracker and unsupervised SC lazebnik2004semi; gupta2023asic; barel2024spacejam. Recent studies have shown that ambiguities caused by symmetric objects are a major source of errors in SC zhang2023telling; mariotti2024improving. One potential way to mitigate these errors is to develop 3D-aware methods. Initially, this has been explored by building correspondence across images by matching points along the surface of objects to 3D meshes zhou2016learning; kulkarni2019canonical; neverova2020continuous; wimmer2023back; dutt2024diffusion, but the requirement for meshes greatly limits the applicability of these approaches. More recently, methods have been proposed to learn 3D shape from image collections monnier2022share; wu2023magicpony; aygun2023saor. However, these methods require solving multiple problems at once, i.e., estimating object shape, deformations, camera pose, and are therefore limited to specific types of object categories and tend to break easily when applied to more complex shapes. Recent advances in monocular depth ranftl2020towards; bhat2023zoedepth; ke2024repurposing; yang2024depth; fu2024geowizard; wang2024moge and geometry prediction wang2024dust3r allows for reliable geometry estimation from a single image, which can be leveraged for imparting 3D-awareness in into SC methods.
Concurrent work wandel2025semalign3d also proposes to geometrically align images in 3D using depth maps to build category prototypes. It is designed around a test-time alignment of the 3D prototype to the test image. This requires knowing the test category beforehand, having build a dedicated prototype for it, having a segmentation mask for the test instance, and solving a computationally expensive alignment problem, limiting its applicability to severely constrained scenarios where the category has been seen during training and throughput it not an issue. In comparison, while our model shares a similar concept of building 3D category prototype, our goal is to design a generalizable SC pipeline by training a correspondence head on top of a backbone. This allows our model to compute correspondence in a single feedforward pass and generalize to new categories all the while not requiring knowledge of the test category or segmentation.
@@ 3.1 Overview
Figure 2: Overview of our approach. We extract segmentation masks and depths maps from training images and backproject object points to produce the posed point clouds $\mathcal{X}^{b}$. We predict dense features with $\Phi$ and match them against our jointly learned sparse category prototype $(\mathcal{P},\mathcal{Z})$ to produce the canonical point clouds $\mathcal{X}^{c}$. The local geometric alignment between the two provides supervision for updating $\Phi$.
[…]
@@
In the standard supervised SC task, we are given a training set, $\{(I^{(1)},\mathcal{K}^{(1)}),\dots,(I^{(N)},\mathcal{K}^{(N)})\}$ where each image $I^{(n)}$ is annotated with a sparse set of semantic keypoints $\mathcal{K}=\{{\bm{k}}_{1},\dots,{\bm{k}}_{\lvert\mathcal{K}\rvert}\}$ where ${\bm{k}}\in\Lambda$. A common strategy, adopted in recent works huang2022learning; cho2021cats; cho2022cats++; luo2023diffusion; taleof2feats; zhang2023telling, is to learn a descriptor function $\Phi(I,\bm{u})$ that produces a local descriptor $\bm{w}$ for each pixel $\bm{u}$ in $I$ such that the descriptors of corresponding keypoints in paired images are close in feature space. While this sparse keypoint supervision helps the model learn semantically meaningful descriptors for the annotated keypoints, it does not guarantee generalization to unlabeled regions of the object.
A promising direction to address this limitation is to incorporate 3D geometry by assigning each pixel a coordinate in an object centric reference frame. Prior works thewlis2017unsupervised; mariotti2024improving explores this idea by projecting object surfaces onto a spherical coordinate system, with each coordinate on the sphere corresponding to a different characteristic point of the object. However, this necessitates inferring both object shape and the viewpoint from a collection of 2D images, a highly ill-posed problem, which requires generating pairs through synthetic warps resulting unrealistic shapes thewlis2017unsupervised or relies on viewpoint supervision which can be challenging to predict automatically. In the next section we show how to combine sparse keypoints annotations with 3D geometry cues to learn dense and semantically consistent descriptors for every pixel in an image. An overview of our approach is shown in Fig.˜2.
@@ 3.2 Canonical Representation Learning
Similar to thewlis2017unsupervised; mariotti2024improving, we aim to learn a 3D canonical representation for each object category, along with a function $\varphi(I,\bm{u})\rightarrow\bm{x}\in\mathbb{R}^{3}$ that maps a pixel $\bm{u}$ in image $I$ to its 3D coordinates in the canonical object-centric coordinate system. Unlike the spherical representations in prior work thewlis2017unsupervised; mariotti2024improving, we do not impose any topological constraints on the canonical representation (e.g., enforcing a spherical surface). We parameterized the canonical representation by a set of 3D keypoints $\mathcal{P}=\{\bm{p}_{1},\dots,\bm{p}_{\lvert\mathcal{K}\rvert}\}$, where each $\bm{p}_{i}$ corresponds to a labeled keypoint ${\bm{k}}_{i}$ in $I$. A crucial aspect of this parametrization is that unlike the labeled image keypoints $\mathcal{K}$, $\mathcal{P}$ is shared across all instances of the category, and is invariant to object instance, viewpoint, and pose, ensuring the canonicity of the representation.
To compute $\mathcal{P}$, we first estimate the 3D coordinates of each keypoint $\bm{k}$} in image $I$ using a monocular depth model $\Psi(I,\bm{k})\rightarrow\mathbb{R}^{+}$ and then backproject it using estimated camera intrinsics $\mathbf{A}\in\mathbb{R}^{3\times 3}$ as follows:
@@
To align $\bar{\mathcal{K}}$ with $\mathcal{P}$, we compute a rigid transformation, comprising rotation $\mathbf{R}\in SO(3)$ and translation $\mathbf{T}\in\mathbb{R}^{3\times 1}$, and scale $s\in\mathbb{R}^{+}$ such that $\mathbf{M}=s[\mathbf{R}|\mathbf{T}]\in\mathbb{R}^{3\times 4}$. We optimize the canonical keypoints $\mathcal{P}$ by minimizing their alignment error across the training set by solving a generalized Procrustes problem:
[…]
We use the Kabsch-Umeyama algorithm kabsch1976solution; umeyama1991least to compute the optimal transformation $\hat{\bm{M}}^{(n)}$ between the canonical and posed coordinates keypoints. To prevent the degenerate solution where $s=0$ leads to a global collapse, we modify the procedure to constrain $s\geq 1$. This ensures objects cannot shrink in size, effectively resizing them to the size of the largest object in the training set. Even though alignments rely on only a sparse subset of visible keypoints per-image, i.e., occluded and out-of-frame points are not considered, we find this sufficient to recover a globally consistent arrangement for $\mathcal{P}$.
@@
with cosine similarity sim(,)($\cdot$$\cdot$) and learned temperature parameter $\tau$. This objective encourages $\Phi$ to produce distinctive and semantically consistent features across object instances for each keypoint $\bm{k}$.